So there are three seats, each one can either be filled by a boy or a girl. How many different combinations are there with repetition?
I know that there are 4:
BBB
BBG
BGG
GGG
However, I don't know the equation to find this... Could you help? Much obliged!

Answer by jim_thompson5910(35256)

There are 2^3 = 8 ways to fill 3 seats with either a boy or a girl. Since order doesn't matter (and something like BBG = BGB and BBG = GBB), we're over-counting and need to divide by 2! = 2*1 = 2 to get

8/2 = 4